# Universal Acceleration of Quantum Relaxation by Reset-Induced Mpemba Effect

### v2

@@ Abstract
Slow relaxation processes spanning widely separated timescales pose fundamental challenges for probing steady-state properties and engineering functional quantum systems, such as quantum heat engines and quantum computing devices. We introduce a universal quantum reset protocol that enables exponential acceleration of relaxation in general Markovian open quantum systems. The protocol induces a quantum Mpemba effect, wherein systems initialized farther from equilibrium can relax faster than those closer to it, and crucially, this acceleration persists even when the slowest decaying Lindbladian modes form complex-conjugate pairs. Unlike previous approaches, which typically target a single mode, our protocol eliminates multiple relaxation modes simultaneously. This framework provides a versatile and experimentally feasible tool for controlling quantum relaxation timescales, with broad implications for quantum thermodynamics, computation, and state preparation.
@@
Researchers have developed various techniques to accelerate relaxation toward both thermal equilibrium [16] and nonequilibrium steady states [17, 18, 11]. Most of these methods, however, are restricted to specific systems and rely on intricate external controls. For instance, some approaches require auxiliary Hamiltonians or full knowledge of the system Hamiltonian, limiting their applicability to simple setups. A more general and elegant strategy is to accelerate relaxation by tailoring the initial state distribution, without the need for continuous external driving. A particularly intriguing example of this approach is inspired by the so-called Mpemba effect [19], a counterintuitive relaxation anomaly in which a hotter system can cool faster than a colder one. This phenomenon implies the existence of an optimal initial state that minimizes the relaxation timescale. Theoretical frameworks for identifying such optimal states have been proposed in both classical [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] and quantum systems [10, 32, 33, 34, 35, 36, 37, 38, 39], see also [40, 41] for recent reviews. However, existing quantum implementations typically impose unrealistic constraints, such as assuming pure initial states or requiring the second-largest eigenvalue of the Lindbladian to be real, which limit their applicability in general settings. Thus, a general and practical method for inducing faster relaxation toward arbitrary steady states in open quantum systems remains elusive. Likewise, a systematic recipe for prolonging the lifetime of metastable states, i.e., for slowing down the relaxation process is also lacking.
In this study, we address these challenges by proposing an experimentally feasible protocol that can both accelerate and decelerate general quantum relaxation processes. The protocol is based on the quantum reset operation [42, 43, 44, 45, 46], which is a key component of many quantum computation tasks [47]. For instance, qubit initialization to the computational ground state $|00...0\rangle$ is typically realized through a reset procedure. Recent experimental advancements have demonstrated the practical implementation of quantum reset, e.g. Aamir et al. have realized an autonomous quantum absorption refrigerator using superconducting circuits, capable of resetting a transmon qubit [48]. Other reset schemes have also been proposed [49, 50, 51, 52]. In our protocol, we introduce a finite-duration reset phase into the open-system dynamics by temporarily coupling the system to a reset channel. We show that this operation can exponentially accelerate relaxation, even when starting from mixed initial states and in the presence of complex decay modes. Remarkably, our method allows for the simultaneous elimination of multiple relaxation modes—a capability absent in previous approaches. Moreover, by applying the protocol selectively, the lifetime of metastable states can be extended, offering a systematic route to suppress relaxation.
@@
[…]
@@
$\mathcal{L}^{\dagger}(O)=i[H,O]+\sum_{i}\left[J_{i}OJ_{i}^{\dagger}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},O\}\right].$
@@
$\rho(t)=e^{t\mathcal{L}}[\rho_{0}]=\rho_{ss}+\sum_{k=2}^{d^{2}}e^{\lambda_{k}t}c_{k}R_{k},$ (3)
where $c_{k}\equiv\text{Tr}(L_{k}^{\dagger}\rho_{0})$. The unique stationary state of the open quantum system $\rho_{ss}$ is given by $\rho_{ss}=\lim_{t\rightarrow\infty}\rho(t)=R_{1}$, assuming $\lambda_{1}=0$ is non-degenerate. In the long-time limit, the relaxation is dominated by the slowest decaying mode, and the deviation from stationarity obeys
$||\rho(t)-\rho_{ss}||\sim\exp\left(|\text{Re}\lambda_{2}|t\right),$
@@
so the characteristic relaxation timescale is given by $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$. For clarity of exposition, we first consider the simple case where the second eigenvalue $\lambda_{2}$ is real and non-degenerate, as in [10]. Later, we will generalize the discussion to include cases where the subleading eigenvalues form complex conjugate pairs.
@@
Reset protocol to accelerate relaxation processes.—Recognizing that the slowest decaying mode governs the relaxation timescale, we propose a quantum reset protocol that suppresses its excitation from a general initial state via a finite-duration reset phase, thereby enabling exponential acceleration of relaxation. Specifically, we let the system evolve under modified dynamics for a finite time $t_{s}$, during which it is stochastically reset to a chosen state $|\delta\rangle$ at random times, with events following a Poisson process of rate $r$. The reset state may also be taken as a mixed state $\rho_{\delta}$. After time $t_{s}$, the reset channel is turned off, and the system resumes evolving under the original Lindbladian $\mathcal{L}$. The modified dynamics during the reset phase is governed by a new Lindbladian $\mathcal{L}_{r}$, defined as
$\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)|\delta\rangle\langle\delta|-r\rho.$ (4)
@@
This can be physically implemented by adding reset jump operators [42] $J_{i}^{r}=\sqrt{r}|\delta\rangle\langle\phi_{i}|$ to the Lindbladian (1), where $|\phi_{i}\rangle$ form an auxiliary complete orthonormal basis. When the reset state is a mixed state $\rho_{\delta}$, the new Lindbladian reads $\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)\rho_{\delta}-r\rho$. Importantly, this protocol modifies the relaxation spectrum without altering the eigenmodes. This is because the term $r\text{Tr}(\rho)|\delta\rangle\langle\delta|$ in the modified Lindbladian contributes only to the stationary mode and vanishes on all traceless eigenmodes $R_{i}$ with $i\geq 2$, which satisfy $\mathrm{Tr}(R_{i})=0$ due to $\mathrm{Tr}(L_{1}^{\dagger}R_{i})=0$ and $L_{1}^{\dagger}=\mathbb{I}$ [cf. (2)]. As a result, the modified Lindbladian acts as
$\mathcal{L}_{r}(R_{i})=(\lambda_{i}-r)R_{i},\ i\in\{2,...,d^{2}\},$ (5)
@@
leading to a uniform spectral shift by $-r$ while preserving the mode structure. Let $\rho^{r}(t):=e^{t\mathcal{L}_{r}}\rho_{0}$ be the state at time $t$ under reset dynamics. Its spectral decomposition reads:
$\rho^{r}(t)=$ $e^{t\mathcal{L}_{r}}\rho_{ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})e^{(\lambda_{k}-r)t}R_{k}$
::=:= $\rho_{ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}R_{k},$ (6)
@@
where $c_{k}^{r}(t)\propto\text{Tr[$L_{k}^{\dagger}$$\rho^{r}(t)$]}$ are time-dependent mode amplitudes. Explicitly, these modified coefficients are given by
@@
where $d_{k}=\langle\delta|L_{k}^{\dagger}|\delta\rangle$. By properly choosing $t_{s}$, we can set $c_{2}^{r}(t_{s})=0$, thereby eliminating the overlap with the slowest decaying mode. Solving $c_{2}^{r}(t_{s})=0$ yields
$t_{s}(r)=\frac{1}{r-\lambda_{2}}\ln\left[1-\frac{c_{2}(r-\lambda_{2})}{r\cdot d_{2}}\right].$ (8)
This condition is always achievable by an appropriate choice of reset state $|\delta\rangle$ ensuring $t_{s}(r)\geq 0$. After the reset phase ($t>t_{s}$), the system returns to original Lindbladian evolution, now without the contribution from the slowest mode. The full evolution reads:
$\rho^{r}(t)=$ $\rho_{ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}\Theta(t_{s}-t)R_{k}$
@@
where $\Theta(x)$ is the Heavside step function. This protocol reduces the relaxation time to
$\tau_{{\rm rel}}\sim t_{s}+\frac{1}{|{\rm Re}(\lambda_{3})|},$ (10)
which is generically shorter than the unperturbed timescale $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$.
Our protocol also enables controlled deceleration of relaxation. By choosing $|\delta\rangle$ and $t_{s}$ to amplify the magnitude of $c_{k}^{r}(t)$, one can extend the lifetime of selected metastable modes. This feature is especially useful in quantum metrology and computation, where long-lived metastable states play a functional role. Our reset protocol thus provides a general framework for controlling relaxation processes in open quantum systems.
@@
$\rho(t)=\rho_{ss}+e^{\lambda_{2}t}c_{2}R_{2}+e^{\lambda_{2}^{\dagger}t}c_{2}^{*}R_{2}^{\dagger}+...,$ (11)
@@
$c_{2}^{r}(t)=\left[c_{2}-\frac{r\cdot d_{2}}{r-\lambda_{2}}\right]e^{-rt}+\frac{r\cdot d_{2}}{r-\lambda_{2}}e^{-\lambda_{2}t}$ ,
$c_{2}^{r,*}(t)=\left[c_{2}^{*}-\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}\right]e^{-rt}+\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}e^{-\lambda_{2}^{*}t}$
@@
because $c_{2}^{r}(t_{s})=0\Rightarrow c_{2}^{r,*}(t_{s})=[c_{2}^{r}(t_{s})]^{*}=0$. This condition can be fulfilled by tuning the reset rate $r$ and the reset duration $t_{s}$, thereby simultaneously eliminating both components of the dominant complex relaxation mode. In practice, it suffices to reduce $|c_{2}^{r}(t_{s})|$ to an exponentially small value ($\text{Re}[c_{2}^{r}(t_{s})]\ll 1,\ \text{Im}[c_{2}^{r}(t_{s})]\ll 1$), which still leads to exponential acceleration of relaxation. We demonstrate this behavior concretely in the following example.
Example.—To illustrate our protocol, we consider a minimal example: a two-level quantum system, which may represent a qubit or a spin-$1/2$ particle. We set the ground state energy $E_{0}=0$ and excited state energy $E_{1}=E$, such that the system Hamiltonian reads
$H=E|1\rangle\langle 1|+\Omega(|0\rangle\langle 1|+|1\rangle\langle 0|)\text{,}$
where $\Omega$ is the intrinsic coherent coupling between the two levels. The system is coupled to a thermal environment through the Lindblad jump operators $J_{0}=|1\rangle\langle 0|=\sigma_{+}$ and $J_{1}=|0\rangle\langle 1|=\sigma_{-}$. The resulting open-system dynamics are governed by the master equation:
$\frac{d\rho(t)}{dt}=$ $-i[H,\rho]+\Gamma_{0}\left(J_{0}\rho J_{1}-\frac{1}{2}\{J_{0}J_{1},\rho\}\right)$
$+\Gamma_{1}\left(J_{1}\rho J_{0}-\frac{1}{2}\{J_{1}J_{0},\rho\}\right),$ (13)
where $\Gamma_{0,1}$ are the transition rates from state $|0\rangle$ to state $|1\rangle$ or converse. Assuming the rates satisfy detailed balance with inverse temperature $\Gamma_{0}e^{-\beta_{\text{env}}E_{0}}=\Gamma_{1}e^{-\beta_{\text{env}}E_{1}}$, i.e., $\Gamma_{0}=\Gamma_{1}e^{-\beta_{\text{env}}E},$ the system evolves to a unique equilibrium state $\rho_{eq}=\frac{e^{-\beta_{\text{env}}H}}{\text{Tr}(e^{-\beta_{\text{env}}H})}$ irrespective of the initial condition. We initialize the system in a general mixed state:
@@
where the off-diagonal terms $ke^{\pm i\phi}$ quantify initial quantum coherence. In the absence of intrinsic transitions ($\Omega=0$), the Hamiltonian is diagonal and the relaxation process reduces to pure decoherence. In this case, all eigenmodes and eigenvalues of the Lindbladian can be computed analytically (see Appendix B). When we apply our reset protocol with the reset state $|\delta\rangle=|0\rangle\langle 0|$, corresponding to the ground state to ensure $t_{s}(r)\geq 0$, the two slowest decaying modes form a complex-conjugate pair:
@@
where $1/r$ characterizes the reset protocol timescale, and $1/|\text{Re}(\lambda_{3})|$ represents the accelerated relaxation timescale with the slowest mode eliminated. Clearly, $\tau_{{\rm rel}}<1/|{\rm Re}(\lambda_{2})|$ for sufficiently large $r$. This structure is general and holds for more complex systems due to analogous spectral modifications introduced by the quantum reset. By appropriately tuning the reset rate, the relaxation timescale can approach $1/|\rm Re(\lambda_{3})|$ arbitrarily closely. Using Eq. (9), the full dynamics under our reset protocol can thus be determined.
Since the condition $c_{2}^{r}(t_{s})=0$ cannot be satisfied exactly in this case—though $|c_{2}^{r}(t_{s})|$ can be made exponentially small by choosing $t_{s}\sim 1/r$—we instead define $t_{s}$ such that a faster decaying mode is suppressed. This pragmatic approach effectively facilitates experimental implementation. Specifically, solving $c_{4}^{r}(t_{s})=0$ gives
@@
where the total transition rate is $k_{\text{tot}}=r+\Gamma_{0}+\Gamma_{1}$. This choice eliminates the fastest decaying relaxation mode, ensuring the slowest mode becomes exponentially negligible for $t_{s}>1/r$. Importantly, alternative selections for $t_{s}$ can also achieve accelerated relaxation, indicating that detailed dynamics knowledge is not essential. Notably, if the initial inverse temperature satisfies $\beta_{0}<\beta_{\text{env}}$, a scenario analogous to “cooling”, the ground state $|0\rangle\langle 0|$ universally serves as an effective reset state regardless of the reset rate $r$. Conversely, when $\beta_{0}>\beta_{\text{env}}$, resetting to the ground state slows relaxation, necessitating instead the excited state $|1\rangle\langle 1|$ as the reset state for acceleration. This example demonstrates that reset-induced acceleration or deceleration of relaxation is readily controllable through appropriate state selection.
To illustrate the effect of the reset protocol, we next present numerical results. We characterize the distance between the transient state $\rho(t)$ and the stationary state $\rho_{ss}$ using the standard quantum trace distance
[…]
where $|A|:=\sqrt{A^{\dagger}A}$. Other measures, such as the L-infinity norm, yield qualitatively similar outcomes. Utilizing this metric, we illustrate the relaxation dynamics by plotting trace distances from $\rho(t)$ to equilibrium $\rho_{\text{eq}}$ over time $t$ for various rate $r$ in Figure 1 (a). The reset scenario employs initial conditions $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$, whereas the non-reset scenario has $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$. With the environment at lower temperature $\beta_{\text{env}}=4.0$, the non-reset system starts closer to equilibrium. Nonetheless, strikingly, the reset protocol substantially accelerates relaxation, manifesting an anomalous effect reminiscent of the quantum Mpemba effect—systems initially farther from stationary state reach stationarity faster. We next plot the critical time $t_{s}$ required to close the reset as a function of the resetting rate $r$ in Figure 1 (b). The results demonstrate that $t_{s}$ decreases monotonically as $r$ increases, with all other parameters held constant. This finding indicates that reset protocols employing larger resetting rates facilitate more rapid relaxation of open quantum systems.

### v3

# Accelerating Quantum Relaxation via Temporary Reset: A Mpemba-Inspired Approach

@@ Abstract
Slow relaxation processes spanning widely separated timescales pose fundamental challenges for probing steady-state properties and engineering functional quantum systems, such as quantum heat engines and quantum computing devices. We introduce a protocol that enables exponential acceleration of relaxation in general Markovian open quantum systems by temporarily coupling the system to a reset channel, inspired by the quantum Mpemba effect. Crucially, this acceleration persists even when the slowest decaying Lindbladian modes form complex-conjugate pairs. Unlike previous approaches, which typically target a single mode, our protocol eliminates multiple relaxation modes simultaneously. Specifically, using a simple depolarizing channel, we show that the relaxation of a dissipative transverse-field Ising chain can be significantly accelerated, underscoring the experimental feasibility of the protocol. This framework provides a versatile tool for controlling quantum relaxation timescales, with broad implications for quantum thermodynamics, computation, and state preparation.
@@
Researchers have developed various techniques to accelerate relaxation toward both thermal equilibrium [16] and nonequilibrium steady states [17, 18, 11]. Most of these methods, however, are restricted to specific systems and rely on intricate external controls. For instance, some approaches require auxiliary Hamiltonians or full knowledge of the system Hamiltonian, limiting their applicability to simple setups. A more general and elegant strategy is to accelerate relaxation by tailoring the initial state distribution, without the need for continuous external driving. A particularly intriguing example of this approach is inspired by the so-called Mpemba effect [19], a counterintuitive relaxation anomaly in which a hotter system can cool faster than a colder one. This phenomenon implies the existence of an optimal initial state that minimizes the relaxation timescale. Theoretical frameworks for identifying such optimal states have been proposed in both classical [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] and quantum systems [10, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], see also [43, 44] for recent reviews. However, existing quantum implementations typically impose strict constraints, such as assuming pure initial states or requiring the second-largest eigenvalue of the Lindbladian to be real. Moreover, careful initial-state design is a delicate task, since it relies on detailed knowledge of the original initial state and the system dynamics, which is typically not known a prior in complex quantum systems. These issues limit their applicability in general settings. Thus, a general and practical method for inducing faster relaxation in open quantum systems, which is robust with respect to different initial states, remains elusive.
In this study, we address these challenges by proposing an experimentally feasible protocol that can both accelerate and decelerate general quantum relaxation processes from any initial state, by resetting the system to a specified (possibly mixed) target state [45, 46, 47]. Here, this type of quantum reset operation can be realized through engineered dissipation, i.e., quantum reservoir engineering [48, 49, 9, 50], and should be distinguished from the conventional reset typically used for qubit initialization in quantum computation [51, 52, 53, 54, 55, 56, 57, 58]. In our protocol, we introduce a finite-duration reset phase into the open-system dynamics by temporarily coupling the system to a reset channel. We show that this protocol can exponentially accelerate relaxation, even when starting from mixed initial states and in the presence of complex decay modes. Remarkably, our method allows for the simultaneous elimination of multiple relaxation modes—a capability absent in previous approaches. Moreover, selective application of the protocol enables the system of interest to remain in metastable states for extended periods, offering a systematic route to suppress relaxation.
@@
Here, $H$ is the Hamiltonian of the system and the jump operators $J_{i}$ describe the dissipative coupling to the environment. The evolution can be analyzed via the spectral decomposition of $\mathcal{L}$, whose eigenvalues are denoted $\lambda_{k}$ and ordered such that $0\geq\text{Re}(\lambda_{k})\geq\text{Re}(\lambda_{k+1})$. Assuming that $\mathcal{L}$ is diagonalizable, let $R_{k}$ and $L_{k}$ denote the corresponding right and left eigenmatrices, satisfying $\mathcal{L}(R_{k})=\lambda_{k}R_{k}$ and $\mathcal{L^{\dagger}}(L_{k})=\lambda_{k}^{\star}L_{k},\ k=1,...,d^{2}$, with biorthogonal normalization
@@
$\mathcal{L}^{\dagger}(O)=i[H,O]+\sum_{i}\left[J_{i}^{\dagger}OJ_{i}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},O\}\right].$
@@
$\rho(t)=e^{t\mathcal{L}}[\rho_{0}]=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}e^{\lambda_{k}t}c_{k}R_{k},$ (3)
@@
where $c_{k}\equiv\text{Tr}(L_{k}^{\dagger}\rho_{0})$. The unique stationary state of the open quantum system $\rho_{\rm ss}$ is given by $\rho_{\rm ss}=\lim_{t\rightarrow\infty}\rho(t)=R_{1}$, assuming $\lambda_{1}=0$ is non-degenerate. In the long-time limit, the relaxation is dominated by the slowest decaying mode, and the deviation from stationarity obeys $||\rho(t)-\rho_{\rm ss}||\sim\exp\left(-|\text{Re}\lambda_{2}|t\right),$ so the characteristic relaxation timescale is given by $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$. For clarity of exposition, we first consider the simple case where the second eigenvalue $\lambda_{2}$ is real and non-degenerate, as in [10]. Later, we will generalize the discussion to include cases where the subleading eigenvalues form complex conjugate pairs.
@@
Reset protocol to accelerate relaxation processes.—Recognizing that the slowest decaying mode governs the relaxation timescale, we propose a quantum reset protocol that suppresses its excitation from a general initial state via a finite-duration reset phase, thereby enabling exponential acceleration of relaxation. Specifically, we let the system evolve under modified dynamics for a finite time $t_{s}$, during which it is stochastically reset to a chosen state $\rho_{\delta}$ at random times, with events following a Poisson process of rate $r$. After time $t_{s}$, the reset channel is turned off, and the system resumes evolving under the original Lindbladian $\mathcal{L}$. That is, the system density matrix $\rho(t)$ evolves under the reset channel for $t\in[0,t_{s}]$ and decouples from the channel for $t>t_{s}$. The modified dynamics during the reset phase is governed by a new Lindbladian $\mathcal{L}_{r}$ [45], defined as
$\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)\rho_{\delta}-r\rho.$ (4)
@@
Expressing $\rho_{\delta}=\sum_{\alpha}p_{\alpha}|\psi_{\alpha}\rangle\langle\psi_{\alpha}|$, $\mathcal{L}_{r}$ is equivalent to introducing jump operators $J_{i}^{r,\alpha}=\sqrt{rp_{\alpha}}|\psi_{\alpha}\rangle\langle\phi_{i}|$ into $\mathcal{L}$, where $|\phi_{i}\rangle$ form an auxiliary complete orthonormal basis (see End Matter). The term $r\text{Tr}(\rho)\rho_{\delta}$ contributes only to the stationary mode and vanishes on all traceless eigenmodes $R_{i}$ with $i\geq 2$, which satisfy $\mathrm{Tr}(R_{i})=0$ due to $\mathrm{Tr}(L_{1}^{\dagger}R_{i})=0$ and $L_{1}^{\dagger}=\mathbb{I}$ [cf. (2)]. Consequently, the modified Lindbladian acts as
$\mathcal{L}_{r}(R_{i})=(\lambda_{i}-r)R_{i},\ i\in\{2,...,d^{2}\}.$ (5)
@@
This induces a uniform spectral shift by $-r$ for all non-stationary eigenmodes, leaving the eigenmatrices ($R_{i}$, $i\geq 2$) themselves unchanged. In contrast, the stationary mode $R_{1}$ retains eigenvalue zero but is modified under reset. Let $\rho^{r}(t):=e^{t\mathcal{L}_{r}}\rho_{0}$ be the state at time $t$ under reset dynamics. Its spectral decomposition reads:
$\rho^{r}(t)$ $=e^{t\mathcal{L}_{r}}\rho_{\rm ss}+\sum_{k=2}^{d^{2}}c_{k}e^{(\lambda_{k}-r)t}R_{k}$
$:=\rho_{\rm ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}R_{k},$ (6)
@@
where $c_{k}^{r}(t)$ are time-dependent mode amplitudes. Explicitly, these modified coefficients are given by (End Matter)
@@
where $d_{k}=\text{Tr}(L_{k}^{\dagger}\rho_{\delta})$. By properly choosing an optimal $t^{*}_{s}$, we can set $c_{2}^{r}(t^{*}_{s})=0$, thereby eliminating the overlap with the slowest decaying mode. Solving $c_{2}^{r}(t_{s})=0$ yields
$t^{*}_{s}(r)=\frac{1}{r-\lambda_{2}}\ln\left[1-\frac{c_{2}(r-\lambda_{2})}{r\cdot d_{2}}\right].$ (8)
By an appropriate choice of $\rho_{\delta}$ such that $c_{2}/d_{2}\leq 0$, $t^{*}_{s}(r)\geq 0$ can be assured. After the reset phase ($t>t_{s}$), the system returns to original Lindbladian evolution, now without the contribution from the slowest mode. The full evolution reads:
$\rho^{r}(t)=$ $\rho_{\rm ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}\Theta(t_{s}-t)R_{k}$
@@
where $\Theta(x)$ is the Heavside step function. This protocol with an optimal $t_{s}=t^{*}_{s}$ reduces the relaxation time to
$\tau_{{\rm rel}}\sim t^{*}_{s}+\frac{1}{|{\rm Re}(\lambda_{3})|},$ (10)
which is generically shorter than the unperturbed timescale $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$. Eq. (9) can be generalized to the dynamics of observables [59], which may be more convenient to measure.
Our protocol also enables controlled deceleration of relaxation. By choosing $\rho_{\delta}$ and $t_{s}$ to amplify the magnitude of $c_{k}^{r}(t)$, one can prolong the time during which the system remains away from the steady state. This capability could be useful in quantum metrology and computation, where long-lived metastable states play a functional role. Our reset protocol thus provides a general framework for controlling relaxation processes in open quantum systems.
@@
$\rho(t)=\rho_{\rm ss}+e^{\lambda_{2}t}c_{2}R_{2}+e^{\lambda_{2}^{\dagger}t}c_{2}^{*}R_{2}^{\dagger}+\mathcal{O}(e^{\lambda_{3}t}),$ (11)
@@
$c_{2}^{r}(t)=\left[c_{2}-\frac{r\cdot d_{2}}{r-\lambda_{2}}\right]e^{-rt}+\frac{r\cdot d_{2}}{r-\lambda_{2}}e^{-\lambda_{2}t}$ , (12a)
$c_{2}^{r,*}(t)=\left[c_{2}^{*}-\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}\right]e^{-rt}+\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}e^{-\lambda_{2}^{*}t}$ (12b)
@@
because $c_{2}^{r}(t_{s})=0\Rightarrow c_{2}^{r,*}(t_{s})=[c_{2}^{r}(t_{s})]^{*}=0$. In some cases, this condition can be exactly satisfied by tuning $r$ and $t_{s}$. Although Eq. (13) does not always admit an exact solution, in practice, it suffices to reduce $|c_{2}^{r}(t_{s})|$ to an exponentially small value ($\text{Re}[c_{2}^{r}(t_{s})]\ll 1,\ \text{Im}[c_{2}^{r}(t_{s})]\ll 1$), which still leads to exponential acceleration of relaxation. Additionally, we show in [59] that, if the condition
${\rm Re}(c_{2}^{*}d_{2})\leq|c_{2}|^{2}$ (14)
is satisfied, $|c_{2}^{r}(t_{s})|<|c_{2}|$ can always be achieved for any positive $t_{s}$ that below a certain threshold. This is a weak condition that be readily satisfied for many choices of $\rho_{\delta}$. It is strictly weaker than the commonly satisfied condition $c_{2}/d_{2}\leq 0$, which, in the case where $\lambda_{2}$ is real, guarantees the existence of a $t_{s}^{*}$ such that $c^{r}_{2}(t_{s}^{*})=0$. Eq. (14) thus provides a sufficient condition for reset-induced Mpemba-type acceleration [20]. It implies that fine-tuning of either $\rho_{\delta}$ or $t_{s}$ is unnecessary for achieving acceleration. We illustrate this behavior concretely in the following examples.
Example 1.—To illustrate our protocol, we first consider a minimal example: a two-level quantum system, which may represent a qubit or a spin-$1/2$ particle. While minimal, this example is crucial in many practical settings, as qubit reset often constitutes the bottleneck in large-scale quantum processes [60, 58]. We set the ground state energy $E_{0}=0$ and excited state energy $E_{1}=E$, such that the system Hamiltonian reads $H=E|1\rangle\langle 1|+\Omega(\sigma^{+}+\sigma^{-})$, where $\Omega$ is the intrinsic coherent coupling between the two levels, $\sigma^{+}=|1\rangle\langle 0|$ and $\sigma^{-}=|0\rangle\langle 1|$. The system is coupled to a thermal environment through the Lindblad jump operators $J_{0}=\sqrt{\gamma_{0}}\sigma^{+}$ and $J_{1}=\sqrt{\gamma_{1}}\sigma_{-}$. Here, $\gamma_{0}$ ($\gamma_{1}$) is the transition rate from state $|0\rangle$ to state $|1\rangle$ (or converse). Assuming the rates satisfy detailed balance with inverse temperature $\gamma_{0}e^{-\beta_{\text{env}}E_{0}}=\gamma_{1}e^{-\beta_{\text{env}}E_{1}}$, i.e., $\gamma_{0}=\gamma_{1}e^{-\beta_{\text{env}}E},$ the system evolves to a unique equilibrium state $\rho_{\rm eq}=\frac{e^{-\beta_{\text{env}}H}}{\text{Tr}(e^{-\beta_{\text{env}}H})}$ irrespective of the initial condition. We initialize the system in a general mixed state:
@@
where the off-diagonal terms $ke^{\pm i\phi}$ quantify initial quantum coherence. In the absence of intrinsic transitions ($\Omega=0$), the Hamiltonian is diagonal and the relaxation process reduces to pure decoherence. In this case, all eigenmodes and eigenvalues of the Lindbladian can be computed analytically [59]. When we apply our reset protocol with the reset state $\rho_{\delta}=|0\rangle\langle 0|$, corresponding to the ground state to ensure $t_{s}(r)\geq 0$, the two slowest decaying modes form a complex-conjugate pair:
@@
where $1/r$ characterizes the reset protocol timescale, and $1/|\text{Re}(\lambda_{3})|$ represents the accelerated relaxation timescale with the slowest mode eliminated. Clearly, $\tau_{{\rm rel}}<1/|{\rm Re}(\lambda_{2})|$ for sufficiently large $r$. By appropriately tuning the reset rate, the relaxation timescale can approach $1/|\rm Re(\lambda_{3})|$ arbitrarily closely. Using Eq. (9), the full dynamics under our reset protocol can thus be determined.
Since the condition $c_{2}^{r}(t_{s})=0$ cannot be satisfied exactly in this case—though $|c_{2}^{r}(t_{s})|$ can be made exponentially small by choosing $t_{s}\sim 1/r$—we instead define $t_{s}$ such that a faster decaying mode is suppressed. Specifically, solving $c_{4}^{r}(t_{s})=0$ gives
@@
where the total transition rate is $k_{\text{tot}}=r+\gamma_{0}+\gamma_{1}$. This choice fully eliminates the fastest decaying relaxation mode. In addition, for a sufficiently large reset rate $r$, the slowest mode’s amplitude becomes exponentially small after the reset phase. Alternative selections for $t_{s}$ can also achieve accelerated relaxation, indicating that detailed dynamics knowledge is not essential. Notably, if the initial inverse temperature satisfies $\beta_{0}<\beta_{\text{env}}$, a scenario analogous to “cooling”, the ground state $|0\rangle\langle 0|$ serves as a useful $\rho_{\delta}$ regardless of the rate $r$. Conversely, when $\beta_{0}>\beta_{\text{env}}$, resetting to the ground state decelerates relaxation, necessitating the choice $\rho_{\delta}=|1\rangle\langle 1|$ for acceleration. This example demonstrates that, by appropriately choosing $\rho_{\delta}$, one can switch between reset-induced acceleration and deceleration of relaxation.
We next present numerical results for illustration. We characterize the distance between the transient state $\rho(t)$ and the stationary state $\rho_{\rm ss}$ using the standard trace distance $D[\rho(t)|\rho_{\rm ss}]:=\text{Tr}|\rho(t)-\rho_{\rm ss}|/2$ where $|A|:=\sqrt{A^{\dagger}A}$. Other measures, such as the $L_{\infty}$ norm $D_{\infty}[\rho(t)|\rho_{\rm ss}]:=\max_{i}|\lambda_{i}|$ [$\lambda_{i}$ is the $i$-th eigenvalue of $\rho(t)-\rho_{\rm ss}$], yield qualitatively similar outcomes. We first plot $D[\rho(t)|\rho_{\rm eq}]$ over $t$ for various rate $r$ in Fig. 1 (a). The parameters for $r>0$ cases are: initial conditions $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$; The $r=0$ scenario has $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$. The environment is at lower temperature $\beta_{\text{env}}=4.0$. Although the $r>0$ cases are initially farther from equilibrium compared to the $r=0$ case, the reset protocol substantially accelerates relaxation so that the former cases reach equilibrium even faster. We next plot $t_{s}$ as a function of $r$ in Fig. 1 (b) and (d), which show that $t_{s}$ decreases monotonically as $r$ increases. Fig. 1 (c) confirms accelerated relaxation for $\Omega=2$, here measured by the $L_{\infty}$ norm to reduce fluctuations in the curves.